OpenStudy (anonymous):

In a formula, z varies inversely as p. If z is 200 when p is 4, find z when p is 20.

zepdrix (zepdrix):

Hey Kenyatta, If z varies inversely with p, the formula involving z and p looks like this,\[\large\rm z=\frac{k}{p}\]Where k is an unknown `constant` of proportionality.

zepdrix (zepdrix):

z=200 when p=4,\[\large\rm z=\frac{k}{p}\qquad\to\qquad 200=\frac{k}{4}\]Solve for k.

zepdrix (zepdrix):

That will be the first step.